OpenStudy (unimatix):

Integral Help (Calc 1)

OpenStudy (unimatix):

\[\int\limits_{}^{} \frac{ 8 }{ 4w^2 + 9 }\]

OpenStudy (unimatix):

The answer I got is different than the given answer, and I want to know if it's the same.

OpenStudy (unimatix):

So I did: \[\frac{ 8 }{ 9 }\int\limits_{}^{}\frac{ 1 }{ \frac{ 4 }{ 9 }w^2+1 }\] v = \[\frac{ 4 }{ 9 } w^2+1\] and I got: \[\frac{ 8 }{ 9 } \ln \frac{ 4 }{ 9 }w^2+1+C\]

OpenStudy (unimatix):

With the 4/9w^2+1 in the brackets of ln. It's showing a trig substitution and I'm wondering if those are the same, or if I've screwed up and can't solve it this way.

OpenStudy (solomonzelman):

that is not the correct (I should say not the ideal, really) u-sub. if you do that sub, where is your dv?

OpenStudy (solomonzelman):

I used "incorrect" because of the absence of expression for dv

OpenStudy (solomonzelman):

have you learned about trig sub?

OpenStudy (unimatix):

Ohhhh. I completely left that out.

OpenStudy (unimatix):

Yeah that makes sense!

OpenStudy (solomonzelman):

Don't do that sub

OpenStudy (solomonzelman):

Even with an expression for dv, don't

OpenStudy (solomonzelman):

Did you learn trigonometric substition?

OpenStudy (unimatix):

Yes I've learned them, sort of. Not really great with them.

OpenStudy (solomonzelman):

Examples: \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{8}{4x^2+9}~dx}\) \(\large\color{black}{\displaystyle\frac{8}{4}\int\limits_{~}^{~}\frac{1}{x^2+\frac{9}{4}}~dx}\) \(\large\color{blue}{\displaystyle x=\frac{3}{2}\tan(\theta)}\)

OpenStudy (solomonzelman):

see why that substitution in blue is the sub you need?

OpenStudy (unimatix):

Not really.

OpenStudy (solomonzelman):

That's ok, I will do an example. \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{1}{x^2+a^2}{~~}dx}\) (where a is a real number constant) This is a general form of a trigonometric substituion where the rule is: \(x=a\tan(\theta)\), and this is what we will do. \(\large\color{black}{\displaystyle x=a\tan\theta }\) \(\large\color{black}{\displaystyle dx=a\sec^2\theta{~}d\theta}\) then substitute for x and for dx accordingly, \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{1}{\left(a\tan\theta\right)^2+a^2}{~~}\left(a\sec^2\theta{~}d\theta\right)}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{a\sec^2\theta}{a^2\tan^2\theta+a^2}{~~}d\theta}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{a\sec^2\theta}{a^2(\tan^2\theta+1)}{~~}d\theta}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{a\sec^2\theta}{a^2(\sec^2\theta)}{~~}d\theta=\int\limits_{~}^{~}\frac{a}{a^2}{~~}d\theta}\) and then integrate and solve for theta

OpenStudy (solomonzelman):

In our case, the \(\color{magenta}{a\tan(\theta)}\) peace is \(\color{blue}{\displaystyle \frac{3}{2}\tan(\theta)}\)

OpenStudy (solomonzelman):

(for the example, recall that a is a constant and so is a²)

OpenStudy (unimatix):

Thank you! I get it now.

OpenStudy (solomonzelman):

Ok, so lets get back to our problem \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{8}{4x^2+9}~dx}\) \(\large\color{black}{\displaystyle\frac{8}{4}\int\limits_{~}^{~}\frac{1}{x^2+\frac{9}{4}}~dx}\) \(\large\color{blue}{\displaystyle x=\frac{3}{2}\tan(\theta)}\)

OpenStudy (solomonzelman):

what will your \(dx\) be equivalent to?